# SCATTER Common Envelope Formalism for Triples

### v1

@@ 3.7 The Functions $\eta$ and $\cal Q$}
One of the most important aspects of the SCATTER formalism is the choice of the functional form of the angular-momentum-transfer efficiency, $\eta$. In DiStefano_2023, $\eta$ was fit to known post-CE binaries. At present, these binaries have provided the only guide to the functional form of $\eta$. We note, however, that the collection and use of more data, even for WD-containing binaries, could influence $\eta$. We may find, for example, that we obtain better fits to the data by employing models in which the two terms in Equation (3) have different values of $\eta$. For binaries containing an NS or BH, the values of the parameters $A$ and $B$, used to compute $\eta$, are likely to be different from those for WDs.
@@
There are two paths toward finding the most physically realistic functional form and parameterization of $\eta$. One is additional observational data. Eventually, the necessary data will be collected and analyzed. For example, GAIA is making it possible to identify a range of binaries in which one component is a compact object (el2021million; shahaf2023triage; yamaguchi2024wide). Thus, we can expect that in the coming years, enough data will exist to constrain $\eta$ for a range of binary types (e.g., those containing BHs).
The second method is through comparison with simulations of individual systems. Although post-CE triples are generally not well constrained, several numerical studies have examined hierarchical triples in which the tertiary donates an envelope that engulfs an inner binary (comerford2020estimating; glanz2021simulations; rosselli2024evolution). A qualitative trend is that the outer orbit can shrink rapidly, while the inner binary can either (i) shrink and sometimes merge or (ii) expand and even become dissociated, depending on system parameters. This behavior points to $\eta$ values for triple systems that can differ from those inferred for binaries. Consider hierarchical triples in which the outer star contributes the CE. For the outer binary, one expects a broadly similar range of $\eta$ (the underlying physics of drag and envelope torques is similar), though possibly with slightly reduced angular-momentum transfer efficiency due to the presence of the inner pair (glanz2021simulations). The $\eta$ formalism can also predict expansion of the inner binary if mass loss from one or both of its components is allowed, or if some systems yield a negative value of $\eta$, for example if the orbits are retrograde.
@@
As an example of how complexities in local physics affect the results of simulations we can follow the results of rosselli2024evolution. They find that the change in inner-binary separation depends sensitively on the dimensionless density-gradient parameter, which we denote $\epsilon_{\rho}$, of the surrounding flow: shallow gradients tend to harden (shrink) the inner binary, while steep gradients tend to soften (expand) it. Hence if we were to follow the results of such simulations there would be an inherent relationship between the density gradient and our $\eta$ parameter.
Thus further data and data analyses, along with more results from simulations, will help us better constrain the functional form and values of $\eta$ in complex systems. For this paper we choose to use the results from the observationally supported WD data (DiStefano_2023). We then explore the uncertainties by varying the values of the parameters $A$ and $B$ that were derived through fits from the data (Section 7).
@@ 4.2.1 Mapping the Pre-CE State to the Post-CE State
The second (solid magenta) curve from the top shows $a_{\mathrm{in}}^{\max}(0)$, the maximum initial radius of the inner orbit consistent with three-body stability while also allowing the outer star to be RL filling. The dashed magenta line at the bottom of the panel shows the corresponding final inner-binary orbital radius, $a_{\mathrm{in}}^{\max}(f)$. We see that, for values of the third star’s core mass as low as $0.2\,M_{\odot}$, the final inner orbital radius is $10^{-2.5}R_{\odot}$. With such a value, the components of the inner binary will merge within the CE, whether they are extended stars or compact objects. For example, even for the maximum value of $M_{3}^{c}$, the time to gravitational merger is only $\sim 17{,}000$ years, also indicative of a likely merger within the CE.
@@ 6.1 Overview
When the triple is not hierarchical, we do not expect it to be dynamically stable at the start of the CE. All things being equal, the population of CE triple-star systems should be dominated by systems that are stable in their pre-CE state, simply because stable states are longer lived. There are, however, situations in which dynamical stability may not be satisfied. These include systems that have recently undergone episodes of stellar and/or binary evolution. In addition, dense stellar environments, such as within globular clusters or near the centers of galaxies, promote interactions between stars that can alter orbits or even produce capture events. Such interactions may yield triples in which the interstellar separations are comparable to each other. Finally, tidal disruptions of a passing star by a binary system may produce the type of system we consider in this section. In this case, the donor is the captured star. Non-hierarchical triple-star states are potentially common.
It is also important to note that the conditions for dynamical stability may not be the only determinants of whether a triple-star state is long lived. This is because non-dynamical physical effects can be significant. For example, tidal forces and/or heavy winds may be in play. Thus, the conditions for stability may be altered.
@@ 6.2 Modeling the Non-Hierarchical Case
The SCATTER formalism relies on a fundamental physical principle: the conservation of angular momentum. To employ angular-momentum conservation, we need to estimate the fraction of the CE with which each star interacts. In DiStefano_2023 and in this paper, the function $\mathcal{Q}_{\delta}(q)$ provides such an estimate. Its functional form is based on the Eggleton function and therefore is related to the RL size, which depends on both gravity and rotation. We expect this definition to be fairly robust for binaries and for those hierarchical triples that can be modelled as if they were independent binaries.
@@
The quantity $Q_{(i,j)}$ represents the fraction of the envelope mass that exchanges angular momentum with the binary formed by stars $i$ and $j$. We apply SCATTER to each binary in the manner of DiStefano_2023. We first compute the fractional change in angular momentum for binary $(i,j)$ by considering the contributions from stars $i$ and $j$ individually:
@@
As in the binary case of DiStefano_2023, the corresponding change in separation is
@@
Now let us take our stellar components to be $1$, 22, and $3$, with Star 1 donating an envelope of mass $M_{1}^{\mathrm{env}}$ and possessing a core mass $M_{1}^{c}$. To simplify the equations, we assume, following DiStefano_2023, that $\eta_{(i,j),i}=\eta_{(i,j),j}\equiv\eta_{ij}$. Under this assumption, we obtain the following expressions for the change in separation of each “binary” in our non-hierarchical triple.
@@ 7.2 Star in the Outer Orbit of a Hierarchical Triple Fills its RL
A primary trend observed in Figure 6 is that orbital shrinkage is maximized when the binary components possess equal masses. Mathematically, this corresponds to the fact that the function $\mathcal{F}(q)$ is maximized at $q=1$ (see Figure 2 of DiStefano_2023). Nevertheless, there are broad ranges of mass ratios, centred around $M_{3}^{c}/M_{\mathrm{bin}}$ and across a broad range of values of $q_{\mathrm{ec}}$, where mergers and/or subsequent episodes of interaction are predicted.
@@
Figure 6: Colour plot illustrating the logarithmic change in outer separation for both inner (top panels) and outer (bottom panels) binaries when the outer star acts as the donor. We generate three sets of plots corresponding to different fitted values of $\eta$ as provided by DiStefano_2023, to cover a representative range of angular-momentum transfer efficiency. These values are then plugged into our SCATTER equations for the inner and outer binary, as described by Equations 7 and 10. The solid white contour represents shrinkage factors beyond $10^{-2}$, signifying conditions where mergers are nearly inevitable. The remaining three contour lines indicate shrinkage factors of $10^{-1.5}$, $10^{-1}$, and unity, respectively.
@@ 8.2 Predictions of Events and Post-CE States
Common-envelope calculations, particularly in the realm of population synthesis, are used to compute a variety of event rates. The event types range from SNe Ia to gravitational mergers of BHs and/or NSs. Correctly predicting event rates is therefore a crucial goal of CE formalisms. Fully incorporating triples into the calculations is crucial to making correct predictions. In fact, triples are capable of producing more events per system than binaries. It is essential to include them because a large fraction of the events and end states we aim to explore emerge from primordial triples, which are common among massive stars.
@@ 8.3.1 Type Ia Supernovae and Other WD Systems
Because the fits to the $\eta$ function are based on post-CE systems containing WDs, SCATTER is ready for simulations designed to compute the rates and times of SNe Ia. With the triple formalism in hand, we can compare the rates and characteristics of events as derived in triple-star systems and in double-star systems.
@@
8.3.2 Gravitational Mergers
In binary scenarios, two BHs must form and then come close enough to each other to merge in a Hubble time. Triple-star systems provide additional opportunities. For example, a pair of BHs may not be close enough to each other to merge within a Hubble time. But if, after the BH binary has formed, there is still a star in an outer orbit, a CE induced by the evolution of that star may either cause the BHs to merge within the CE or else to be brought close enough during the CE that they will merge in a Hubble time. Furthermore, the remnant of the outer star could merge with the merged remnant of the inner star.
@@ Acknowledgements
The authors thank Yan Gao for discussions. RD and CK also thank Matthias Kruckow and Patrick Neunteufel for discussions. R.D. acknowledges support from the National Science Foundation, through NSF AST-2009520. C.K. acknowledges funding from the UK Science and Technology Facility Council through grant ST/Y001443/1.

### v2

@@ 3.7 The Functions $\eta$ and $\cal Q$}
One of the most important aspects of the SCATTER formalism is the choice of the functional form of the angular-momentum-transfer efficiency, $\eta$. In Di Stefano et al. (2023), $\eta$ was fit to known post-CE binaries. At present, these binaries have provided the only guide to the functional form of $\eta$. We note, however, that the collection and use of more data, even for WD-containing binaries, could influence $\eta$. We may find, for example, that we obtain better fits to the data by employing models in which the two terms in Equation (3) have different values of $\eta$. For binaries containing an NS or BH, the values of the parameters $A$ and $B$, used to compute $\eta$, are likely to be different from those for WDs.
@@
There are two paths toward finding the most physically realistic functional form and parameterization of $\eta$. One is additional observational data. Eventually, the necessary data will be collected and analyzed. For example, GAIA is making it possible to identify a range of binaries in which one component is a compact object (El-Badry et al., 2021; Shahaf et al., 2023; Yamaguchi et al., 2024). Thus, we can expect that in the coming years, enough data will exist to constrain $\eta$ for a range of binary types (e.g., those containing BHs).
The second method is through comparison with simulations of individual systems. Although post-CE triples are generally not well constrained, several numerical studies have examined hierarchical triples in which the tertiary donates an envelope that engulfs an inner binary (Comerford & Izzard, 2020; Glanz & Perets, 2021; Rosselli-Calderon et al., 2024). A qualitative trend is that the outer orbit can shrink rapidly, while the inner binary can either (i) shrink and sometimes merge or (ii) expand and even become dissociated, depending on system parameters. This behavior points to $\eta$ values for triple systems that can differ from those inferred for binaries. Consider hierarchical triples in which the outer star contributes the CE. For the outer binary, one expects a broadly similar range of $\eta$ (the underlying physics of drag and envelope torques is similar), though possibly with slightly reduced angular-momentum transfer efficiency due to the presence of the inner pair (Glanz & Perets, 2021). The $\eta$ formalism can also predict expansion of the inner binary if mass loss from one or both of its components is allowed, or if some systems yield a negative value of $\eta$, for example if the orbits are retrograde.
@@
As an example of how complexities in local physics affect the results of simulations we can follow the results of Rosselli-Calderon et al. (2024). They find that the change in inner-binary separation depends sensitively on the dimensionless density-gradient parameter, which we denote $\epsilon_{\rho}$, of the surrounding flow: shallow gradients tend to harden (shrink) the inner binary, while steep gradients tend to soften (expand) it. Hence if we were to follow the results of such simulations there would be an inherent relationship between the density gradient and our $\eta$ parameter.
Thus further data and data analyses, along with more results from simulations, will help us better constrain the functional form and values of $\eta$ in complex systems. For this paper we choose to use the results from the observationally supported WD data (Di Stefano et al., 2023). We then explore the uncertainties by varying the values of the parameters $A$ and $B$ that were derived through fits from the data (Section 7).
@@ 4.2.1 Mapping the Pre-CE State to the Post-CE State
The second (solid magenta) curve from the top shows $a_{\mathrm{in}}^{\max}(0)$, the maximum initial radius of the inner orbit consistent with three-body stability while also allowing the outer star to be RL filling. The dashed magenta line at the bottom of the panel shows the corresponding final inner-binary orbital radius, $a_{\mathrm{in}}^{\max}(f)$. We see that, for values of the third star’s core mass as low as $0.2\,M_{\odot}$, the final inner orbital radius is $10^{-2.5}R_{\odot}$. With such a value, the components of the inner binary will merge within the CE, whether they are extended stars or compact objects. For example, even for the maximum value of $M_{3}^{c}$, the time to gravitational-wave merger is only $\sim$17,00017{,}000 years, also indicative of a likely merger within the CE.
@@ 6.1 Overview
When the triple is not hierarchical, we do not expect it to be dynamically stable at the start of the CE. All things being equal, the population of CE triple-star systems should be dominated by systems that are stable in their pre-CE state, simply because stable states are longer lived. There are, however, situations in which dynamical stability may not be satisfied. These include systems that have recently undergone episodes of stellar and/or binary evolution. In addition, dense stellar environments, such as within globular clusters or near the centers of galaxies, promote interactions between stars that can alter orbits or even produce capture events. Such interactions may yield triples in which the interstellar separations are comparable to each other. Finally, tidal disruptions of a passing star by a binary system may produce the type of system we consider in this section. It is also important to note that the conditions for dynamical stability may not be the only determinants of whether a triple-star state is long lived. This is because non-dynamical physical effects can be significant. For example, tidal forces and/or heavy winds may be in play. Thus, the conditions for stability may be altered. Although we expect that non-hierarchical triples form at most a small fraction of all triples experiencing a CE, we consider them for the sake of completeness.
@@ 6.2 Modeling the Non-Hierarchical Case
The SCATTER formalism relies on a fundamental physical principle: the conservation of angular momentum. To employ angular-momentum conservation, we need to estimate the fraction of the CE with which each star interacts. In Di Stefano et al. (2023) and in this paper, the function $\mathcal{Q}_{\delta}(q)$ provides such an estimate. Its functional form is based on the Eggleton function and therefore is related to the RL size, which depends on both gravity and rotation. We expect this definition to be fairly robust for binaries and for those hierarchical triples that can be modelled as if they were independent binaries.
@@
The quantity $Q_{(i,j)}$ represents the fraction of the envelope mass that exchanges angular momentum with the binary formed by stars $i$ and $j$. We apply SCATTER to each binary in the manner of Di Stefano et al. (2023). We first compute the fractional change in angular momentum for binary $(i,j)$ by considering the contributions from stars $i$ and $j$ individually:
@@
As in the binary case of Di Stefano et al. (2023), the corresponding change in separation is
@@
Now let us take our stellar components to be $1$, 22, and $3$, with Star 1 donating an envelope of mass $M_{1}^{\mathrm{env}}$ and possessing a core mass $M_{1}^{c}$. To simplify the equations, we assume, following Di Stefano et al. (2023), that $\eta_{(i,j),i}=\eta_{(i,j),j}\equiv\eta_{ij}$. Under this assumption, we obtain the following expressions for the change in separation of each “binary” in our non-hierarchical triple.
@@ 7.2 Star in the Outer Orbit of a Hierarchical Triple Fills its RL
A primary trend observed in Figure 6 is that orbital shrinkage is maximized when the binary components possess equal masses. Mathematically, this corresponds to the fact that the function $\mathcal{F}(q)$ is maximized at $q=1$ (see Figure 2 of Di Stefano et al. 2023). Nevertheless, there are broad ranges of mass ratios, centred around $M_{3}^{c}/M_{\mathrm{bin}}$ and across a broad range of values of $q_{\mathrm{ec}}$, where mergers and/or subsequent episodes of interaction are predicted.
@@
Figure 6: Colour plot illustrating the logarithmic change in outer separation for both inner (top panels) and outer (bottom panels) binaries when the outer star acts as the donor. We generate three sets of plots corresponding to different fitted values of $\eta$ as provided by Di Stefano et al. (2023), to cover a representative range of angular-momentum transfer efficiency. These values are then plugged into our SCATTER equations for the inner and outer binary, as described by Equations 7 and 10. The solid white contour represents shrinkage factors beyond $10^{-2}$, signifying conditions where mergers are nearly inevitable. The remaining three contour lines indicate shrinkage factors of $10^{-1.5}$, $10^{-1}$, and unity, respectively.
@@ 8.2 Predictions of Events and Post-CE States
Common-envelope calculations, particularly in the realm of population synthesis, are used to compute a variety of event rates. The event types range from SNe Ia to gravitational-wave mergers of BHs and/or NSs. Correctly predicting event rates is therefore a crucial goal of CE formalisms. Fully incorporating triples into the calculations is crucial to making correct predictions. In fact, triples are capable of producing more events per system than binaries. It is essential to include them because a large fraction of the events and end states we aim to explore emerge from primordial triples, which are common among massive stars.
@@ 8.3.1 Type Ia Supernovae and Other WD Systems
Because the fits to the $\eta$ function are based on post-CE systems containing WDs, SCATTER is ready for simulations designed to compute the rates and delay-times of SNe Ia. With the triple formalism in hand, we can compare the rates and characteristics of events as derived in triple-star systems and in double-star systems.
@@
8.3.2 Gravitational-Wave Mergers
In binary scenarios, two BHs must form and then come close enough to each other to merge in a Hubble time. Triple-star systems provide additional opportunities. For example, a pair of BHs may not be close enough to each other to merge within a Hubble time. But if, after the BH binary has formed, there is still a star in an outer orbit, a CE induced by the evolution of that star may either cause the BHs to merge within the CE or else to be brought close enough during the CE that they will merge in a Hubble time. Furthermore, the remnant of the outer star could merge with the merged remnant of the inner star, as considered in Stegmann et al. (2022).
@@ 8.4 Further Development of the Formalism
  • We have considered circular orbits. While the population of triples includes systems in which an inner or outer binary may be eccentric, we do not explicitly include eccentricity in the CE calculations described in this paper. Note, however, that eccentricity could have played a role during the pre-CE evolution. It can also play a role once the CE has been dispersed. It will be useful to extend the SCATTER formalism to be to take into account the role of eccentricity during the CE itself; it would circularize the system with an initial eccentricity assumed. This will allow explicit consideration of a number of interesting cases. For example, if the outer orbit is eccentric, the star in the outer orbit may plunge into a CE that encompasses the inner two stars.
@@ Acknowledgements
The authors thank Yan Gao for discussions. RD and CK also thank Matthias Kruckow and Patrick Neunteufel for discussions. R.D. acknowledges support from the National Science Foundation, through NSF AST-2009520. C.K. acknowledges funding from the UK Science and Technology Facility Council through grant ST/Y001443/1. We thank the referee, Jakob Stegmann, for helpful comments.